Find dy/dx of the function xy2 + e6x y3 = cos(x2 + 2) at the point (1,1) by using implicit differentiation. 


1
Expert's answer
2022-02-07T17:51:59-0500

Using implicit differentiation we get

"y^2+2xyy'+6e^{6x}y^3+e^{6x}3y^2y'=-\sin(x^2+2)2x."

It follows that at the point "(1,1)" we have

"1+2y'+6e^{6}+3e^{6}y'=-2\sin(3)."

Therefore,

"(2+3e^{6})y'=-1-6e^{6}-2\sin(3),"

and we conclude that

"\frac{dy}{dx}=y'=-\frac{1+6e^{6}+2\sin(3)}{2+3e^{6}}."